OpenStudy (mathmath333):

Find the probablity that you get a total of more than 3 in a throw of 2 dice.

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ Find the probablity that you get a total of more than }\hspace{.33em}\\~\\ & \normalsize \text{ 3 in a throw of 2 dice .}\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (badhi):

more than 3 -> 6x4 (1,2,3,4,5,6) and (3,4,5,6) all the possibilites -> 6x6 (1,2,3,4,5,6) and (1,2,3,4,5,6)

OpenStudy (mathmath333):

is it 24/36=2/3

OpenStudy (anonymous):

you could just do \[\Pr(T>3)=1-\Pr(T \le3)\]

OpenStudy (anonymous):

would be less ways that way

OpenStudy (anonymous):

and you know there are 36 combinations in total for rolling any number.

OpenStudy (badhi):

yes chris is correct.. btw i forgot about (3,1), (4,1),... (3,2), (4,2),... combinations I think its better to follow chris 's way

OpenStudy (anonymous):

1,1 2,1 2,1 are the only combinations less than or equal to 3

OpenStudy (mathmath333):

is it 11/12

OpenStudy (anonymous):

\[\Pr(X \le3)=\frac{ 1 }{ 36 }+\frac{ 2 }{ 36 }\]

OpenStudy (anonymous):

ye

OpenStudy (anonymous):

looks good

OpenStudy (mathmath333):

11/12 is correct

OpenStudy (anonymous):

awesome